OpenStudy (anonymous):

Solve sin(x)(sin(x)-1) = 0

OpenStudy (anonymous):

@hartnn I got sin(x) = 1/2. None of the choices show 30 degrees.

hartnn (hartnn):

ab=0 gives a=0 or b= 0 what about sin(x)(sin(x)-1) = 0 ? sin x =1/2 ? O.o how ?

OpenStudy (anonymous):

I moved the 1 over and combined the sins to make 2sin....

OpenStudy (anonymous):